cotA-tanA=2. the principal value of A is?
  1. 45    b. 90   c. 22.5   d.135

Grade:11

1 Answers

Vikas TU
7 years ago
Write, cotA-tanA=2.  as:
1/tanA – tanA = 2
let tanA = t
1/t – t = 2
1 – t^2 = 2t
or
t^2 + 2t – 1 = 0
D = 4 + 4 > 0 
t = (-2 + 2root2)/2 = -1 + roott2
and = -1 – root2